# A Tactile Void

### v1

@@ Abstract
where $\alpha=2a/3\sqrt{1/(3\pi RLE)}$ and $\beta=z^{2}\pi LE/(3R)$ are two constants. The experimental measurements of $u_{r}$ are well adjusted by a $\cos{(2\theta+\varphi)}$ function (see Fig. 2D), where $\varphi\sim\pi/8$ is a phase shift required to adjust the data, likely due to the fact that the cavity is not perfectly located in the contact symmetry plane (see Fig. 1 A4). The amplitude of the $u_{r}$ modulations under normal load, $U_{r}^{N}$, are extracted from the standard deviation of the $u_{r}(\theta)$ curves, and are plotted against the normal force on Fig.2E. The data are reasonably well fitted with Eq.2, with $\alpha=11\pm 2\mu$m/$N^{1/2}$ and $\beta=1.9\pm 1$N, the order of magnitude of which compares to the theoretical values $\alpha\approx 24\mu$m/$N^{1/2}$ and $\beta=0.8$ N.
@@
Altogether, this work propose a mechanical framework to predict mechanoreceptor deformations upon stereotypical tactile tasks. Is this simplified model applicable to natural mechanoreceptors ? The typical membrane tensions $\gamma$ triggering the response of biological mechanoreceptors lie around the mN/m scale. Taking a typical value of $E=$ 100 kPa for the Young’s modulus of the human finger [24] one obtains an elastocapillary length [25] $l_{ec}=\gamma/E\sim 10$ nm, extremely small with respect to the typical mechanoreceptor diameter (tens to hundreds of $\mu$m). As a consequence, elastic stresses are dominant to predict the mechanoreceptor deformations and one can safely neglect interfacial contributions. Using this elastic model under static normal load, we can estimate the order of magnitude of the strain at the cavity wall as $\epsilon\approx\ln(1+U_{r}/a)\approx 0.02$. To this strain corresponds a stress $\sigma_{\theta\theta}\sim Eu_{r}^{N}/a\sim\sigma_{0}$ of the order of 10 kPa. For biological mechanoreceptors, to what extent does the membrane tension $\Delta\gamma$ increases upon contact stress ? It can be estimated from the typical area expansion modulus $K\sim$ 100 mN/m of a cell membrane [26, 27]: $\Delta\gamma=K\frac{\Delta\mathcal{A}}{\mathcal{A}}$, with $\mathcal{A}$ the membrane surface. One can estimate $\Delta\gamma\sim U_{r}^{2}/a^{2}\sim K(\frac{\sigma_{0}}{E})^{2}$. With the aforementioned value of $E$ one finds $\Delta\gamma\sim$ 1 mN/m, a value in the range of opening threshold tensions of mechanosensitive proteins[6, 10].
@@
Secondly, we notice from Eq. 2 that interestingly the hoop stress is anisotropic, being compressive at the cavity equator and extensional at the poles. This observation is known in the context of geophysical studies [28, 29] to explain the localization of the fracture plane in gas cavities embedded in rocks. Such a stress anisotropy has never been mentioned in the context of touch mechanoreceptors. Back to the biological problem, this suggests that depending on their angular position on the membrane, the mechanosensitive protein pores will not have the same probability of opening. One may further wonder whether the density of mechanosensitive proteins in the mechanoreceptor membrane is isotropic, or whether Evolution has favored an anisotropic distribution (at the poles and not at the equator) to maximize tactile sensitivity. A second - more likely - hypothesis would be that the angular distribution of protein density changes in response to the local membrane curvature by a negative curvotaxis-like process, as reported for many different membrane proteins [30, 31, 32, 33].
@@
Considering texture discriminations, our findings suggest that shape fluctuations, and as a consequence membrane stress fluctuations, allows for roughness discrimination of probed surfaces. This study calls for mechanical models of cavities under friction forces to assess this experimental finding. On the biological side, it brings a mechanical reason of being for Fast Adapting mechanoreceptors, whose neural response is sensitive to stress fluctuations.

### v2

@@ III Results
where $\alpha=2a/3\sqrt{1/(3\pi RLE)}$ and $\beta=z^{2}\pi LE/(3R)$ are two constants. The experimental measurements of $u_{r}$ are well adjusted by a $\cos{(2\theta+\varphi)}$ function (see Fig. 2D), where $\varphi\sim\pi/8$ is a phase shift required to adjust the data,due to the fact that the cavity is not perfectly located in the contact symmetry plane (see Fig. 1 A4 and dedicated measurements in the Supplementary Materials). The amplitude of the $u_{r}$ modulations under normal load, $U_{r}^{N}$, are extracted from the standard deviation of the $u_{r}(\theta)$ curves, and are plotted against the normal force on Fig.2E. The data are reasonably well fitted with Eq.2, with $\alpha=11\pm 2\mu$m/$N^{1/2}$ and $\beta=1.9\pm 1$N, the order of magnitude of which compares to the theoretical values $\alpha\approx 24\mu$m/$N^{1/2}$ and $\beta=0.8$ N. The cavity’s response sensitivity $\chi$ to normal indentation can be calculated from Eq.2, it writes :
$\chi=\frac{1}{a}\left(\frac{\partial U_{r}^{N}}{\partial F_{n}}\right)_{F_{n}\rightarrow 0}=\frac{\alpha}{a\sqrt{\beta}}=\frac{2}{3\pi}\frac{1}{ELz}$ (3)
The sensitivity is expected to scale as $1/z$ and to diverge as the cavity’s location approaches the cylinder contact plane.
@@ IV Discussion
Altogether, this work propose a mechanical framework to predict mechanoreceptor deformations upon stereotypical tactile tasks. Is this simplified model applicable to natural mechanoreceptors ? The typical membrane tensions $\gamma$ triggering the response of biological mechanoreceptors lie around the mN/m scale. Taking a typical value of $E=$ 100 kPa for the Young’s modulus of the human finger [24] one obtains an elastocapillary length [25] $l_{ec}=\gamma/E\sim 10$ nm, extremely small with respect to the typical mechanoreceptor diameter (tens to hundreds of $\mu$m). As a consequence, elastic stresses are dominant to predict the mechanoreceptor deformations.
Using this elastic model, we found that the static contact sensitivity $\chi$ increases and diverges as the cavity approaches the cylinder apex (Eq. 3). However, from a biological perspective, a balance is likely required between tactile sensitivity and the structural integrity of mechanoreceptors during tactile exploration. Merkel corpuscles, specialized touch mechanoreceptors responsible for detecting static contact, are found approximately 1 mm beneath the skin surface, within dermal papillae (see, for example, [1]). These corpuscles are situated just below the epidermis, which may possibly act as a protective barrier.
Figure 4: See also [26] A) Sketch of a pseudo-vesicle trapped in an agar gel. B) Fluorescence image of the pseudo vesicle, whose internal aqueous phase is marked with a green fluorophore and the oil phase in red (scale bar 200 $\mu$m). C) Sketch of the plane/plane deformation setup. D) Fluorescence images of the pseudo-vesicle upon deformation. The scale bar is 400 $\mu$m long. E) Radial deformation $U_{R}^{N}$ as a function of the normal force, for gas cavities (red discs) a liquid cavities (green triangles) and pseudo-vesicle (blue stars). F) Rescaled $U_{r}^{N}/a$ as a function of rescaled stress $\sigma/E$.
In addition, natural mechanoreceptors embed a complex biological aqueous content within a plasma membrane, made of a lipid bilayer in which various membrane proteins are inserted. To add an extra layer of complexity, the external extracellular matrix is also viscoelastic. To what extent do these properties modify the mechanical response of natural mechanoreceptors ? To take a step closer to replicating natural mechanoreceptors, we previously developed lipid vesicle-like systems embedded in a hydrogel. In a recent paper [26], we demonstrated how to create and trap a lipid pseudo-vesicle within an aqueous agar gel. The term pseudo-vesicle refers to a lipid vesicle with an oily cap on top (Fig. 4A and B).
The agar gel can be indented using a piston with a cross-sectional area $S$ (Fig. 4C and D), and its deformation $u_{r}^{N}$ can be measured as a function of the normal load in a plane-plane contact. In the Supplementary Materials of [26], we showed that the deformation of the pseudo-vesicle closely approximates that of a gas cavity in a PDMS elastomer, as studied in the present work.
Notably, when the deformation $U_{r}^{N}$ is rescaled by the radius $a$ of the pseudo-vesicle (or cavity), and the applied compressive stress $\sigma=F_{n}/S$ is normalized by the Young’s modulus $E$, the curves for different systems collapse onto a single trend (Fig. 4E and F).
These results constitute additional clues to support our modeling of mechanoreceptor deformations using the approximation of a gas cavity, as a reasonable first approach for predicting their behavior in typical tactile tasks.
Using this elastic model under static normal load, we can estimate the order of magnitude of the strain at the cavity wall as $\epsilon\approx\ln(1+U_{r}/a)\approx 0.02$. To this strain corresponds a stress $\sigma_{\theta\theta}\sim Eu_{r}^{N}/a\sim\sigma_{0}$ of the order of 10 kPa. For biological mechanoreceptors, to what extent does the membrane tension $\Delta\gamma$ increases upon contact stress ? It can be estimated from the typical area expansion modulus $K\sim$ 100 mN/m of a cell membrane [27, 28]: $\Delta\gamma=K\frac{\Delta\mathcal{A}}{\mathcal{A}}$, with $\mathcal{A}$ the membrane surface. One can estimate $\Delta\gamma\sim U_{r}^{2}/a^{2}\sim K(\frac{\sigma_{0}}{E})^{2}$. With the aforementioned value of $E$ one finds $\Delta\gamma\sim$ 1 mN/m, a value in the range of opening threshold tensions of mechanosensitive proteins[6, 10].
@@
Secondly, we notice from Eq. 2 that the anistropic experimental radial displacement $U_{r}^{N}$ yields to a theoretically anisotropic hoop stress $\sigma_{\theta\theta}$, being compressive at the cavity equator and extensional at the poles. This observation is known in the context of geophysical studies [29, 30] to explain the localization of the fracture plane in gas cavities embedded in rocks. Such a stress anisotropy has never been mentioned in the context of touch mechanoreceptors. Back to the biological problem, this suggests that depending on their angular position on the membrane, the mechanosensitive protein pores will not have the same probability of opening. One may further wonder whether the density of mechanosensitive proteins in the mechanoreceptor membrane is isotropic, or whether Evolution has favored an anisotropic distribution (at the poles and not at the equator) to maximize tactile sensitivity. A second - more likely - hypothesis would be that the angular distribution of protein density changes in response to the local membrane curvature by a negative curvotaxis-like process, as reported for many different membrane proteins [31, 32, 33, 34].
@@
Considering texture discriminations, our findings suggest that shape fluctuations, and as a consequence membrane stress fluctuations, allows for roughness discrimination of probed surfaces. This study calls for mechanical models of cavities under friction forces to assess this experimental finding. A first step in this direction is to measure and model the cavity’s response to the passage of an elementary texture (a single defect), as we outline in the Supplementary Materials. On the biological side, our results brings a mechanical reason of being for Fast Adapting mechanoreceptors, whose neural response is sensitive to stress fluctuations.